# envname-P envname#1

### v1

@@ Abstract.
Reinforcement Learning (RL)-Based Recommender Systems (RSs) are increasingly recognized for their ability to improve long-term user engagement. Yet, the field grapples with challenges such as the absence of accessible frameworks, inconsistent evaluation standards, and the complexity of replicating prior work. Addressing these obstacles, we present EasyRL4Rec, a user-friendly and efficient library tailored for RL-based RSs. EasyRL4Rec features lightweight, diverse RL environments built on five widely-used public datasets, and is equipped with comprehensive core modules that offer rich options to ease the development of models. It establishes consistent evaluation criteria with a focus on long-term impacts and introduces customized solutions for state modeling and action representation tailored to recommender systems. Additionally, we share valuable insights gained from extensive experiments with current methods. EasyRL4Rec aims to facilitate the model development and experimental process in the domain of RL-based RSs. The library is openly accessible at https://github.com/chongminggao/EasyRL4Rec.
Recommender Systems, Reinforcement Learning, Code Library
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems (RSs) are increasingly becoming integral to various domains, such as e-commerce, social media, and online streaming services. Traditional RSs usually rely on supervised learning methods like collaborative filtering (Schafer et al., 2007) to learn users’ interests and predict their immediate responses. However, these methods often fail to capture the long-term effects, leading to potential issues like Feedback Loops (Mansoury et al., 2020), Filter Bubbles (Flaxman et al., 2016), and other undesirable outcomes.
Reinforcement Learning (RL)-Based Recommender Systems have gained rising attention due to their ability to optimize long-term user engagement. In this setting, the entire recommendation process is perceived as a multi-step decision-making problem. An agent (recommender) interacts with the environment (users) and receives feedback on a specific action. The main goal of RL-based RSs is to learn an optimal recommendation policy that maximizes cumulative rewards (users’ long-term engagement) through trial and error. Currently, RL-based RSs have been extensively applied in scenarios such as short video recommendations (Cai et al., 2023b; Liu et al., 2023; Cai et al., 2023a; Xue et al., 2023), e-commerce recommendations (Rohde et al., 2018; Zhao et al., 2018a), and beyond.
With increasing attention to RL-based RS, research in this field is confronted with the following three challenges.
  • Absence of an easy-to-use framework. Currently, there is a lack of easy-to-use frameworks or libraries for the study of RL-based RS. Research for RL-based RS requires interactive environment building. However, existing implementations of environments neither use mass datasets nor simulated data generated by simulators, which is inconvenient and hard to follow for most researchers. In addition, it is difficult to directly apply RL libraries (e.g. rllib (Liang et al., 2018)) to recommendation scenarios due to the lack of recommendation environment construction, and state modeling.
@@
To tackle the above issues, we implement a comprehensive and easy-to-use code library for RL-based RSs, named EasyRL4Rec. It provides a user-friendly framework with core modules with rich choices and a unified training and evaluating process, aiming to simplify the model development and experimental process in the domain of RL-based RS. The library is structured around four core modules: Environment, Policy, StateTracker, and Collector. These modules cater to different stages of the reinforcement learning interaction process. Environment, built from lightweight static datasets, provides feedback on upcoming actions. The Policy module selects the optimal action based on the current state, which is encoded by StateTracker. Collector bridges the interactions between Environment and Policy. To address the difficulty of obtaining user states from environments, we implement multiple StateTrackers for state modeling, encompassing today’s popular methods in sequential modeling (Hidasi et al., 2015; Tang and Wang, 2018; Kang and McAuley, 2018; Yuan et al., 2019). Since items in recommendation systems are discrete, EasyRL4Rec includes a mechanism to convert continuous actions to discrete items, allowing for continuous action-based policies.
Moreover, EasyRL4Rec provides a unified training and evaluation procedure by Trainer and Evaluator executors. Evaluator supports three modes, allowing for the removal of recommended items and a quit mechanism. Trainer offers two training paradigms: learning directly from offline logs or training with a pre-trained user model. With this unified and modular framework, we conduct comprehensive experiments on classic RL models and several recent work (Xin et al., 2020; Gao et al., 2023). The experimental results uncovered several intriguing insights, such as the Preference Overestimation issue in offline RL.
The main contributions and key features of this library can be summarized as follows:
  • Easy-to-use Framework. EasyRL4Rec provides an easy-to-use framework for RL-based Recommendations. We construct lightweight RL environments based on five public datasets encompassing diverse domains, which are easy to follow for researchers. Moreover, the design of core modules with rich options reduces the complexity of developing a new model.
@@
  • Tailored Designs for Recommendation Scenarios. In response to challenges when applying RL algorithms in practical recommender systems, we have developed customizable modules for state modeling and action representation. We implement multiple StateTrackers for state modeling. EasyRL4Rec includes a mechanism to convert continuous actions to discrete items, supporting continuous action-based policies.
  • Insightful Experiments for RL-based RSs. With EasyRL4Rec, we conduct comprehensive experiments to compare the performance of classic RL models and some recent work. The experimental results uncovered several intriguing insights, such as the Preference Overestimation issue in offline RL.
@@ 2.1. Development of RL-Based RSs
Reinforcement Learning (RL), a branch of machine learning, focuses on agents learning the ability of decision-making through environmental feedback [^1] RL in this work specifically refers to deep reinforcement learning.. Classical RL algorithms can be generally divided into three groups: value-based (e.g. DQN (Mnih et al., 2013)), policy gradient-based (e.g. REINFORCE (Williams, 1992)), and actor-critic-based (e.g. PPO (Schulman et al., 2017)). Recently, Reinforcement Learning (RL)-Based Recommender Systems (RSs) have gained considerable attention due to their ability to model recommendation as a multi-step decision-making process and enhance the long-term benefits.
Numerous studies have investigated the application of RL in recommender systems.  Shani et al. (2005) first formulated the recommendation process as a Markov Decision Process (MDP) and utilized model-based RL methods. Zhao et al. (2018b) adapt a DQN architecture to incorporate positive and negative feedback from users.  Zheng et al. (2018) firstly applies the Dueling DQN algorithm to news recommendation. Chen et al. (2019, 2022) extend REINFORCE (Williams, 1992) and off-policy actor-critic algorithm to recommendation.  Xin et al. (2020) proposed to utilize self-supervision signals to empower RL-based RSs. Unlike most methods based on discrete actions,  Cai et al. (2023a); Liu et al. (2023); Xue et al. (2023) investigate the applications of continuous action-based RL and improvements of long-term user engagement in short-video scenarios. In addition,  Ren et al. (2023) focus on effective state representations, while recent work (Gao et al., 2022c, 2023) addresses specific issues (e.g. Filter Bubbles, Matthew Effect) when applying RL in RSs.
@@ 2.2. Resources for RL-Based RSs
Existing resources for RL-Based RSs can be categorized as two groups: Simulators & Datasets, and Frameworks & Libraries. In Table  1, we summarize the characteristics of existing resources and our EasyRL4Rec regarding modules and experimental procedures.
@@ 2.2.1. Simulators and Datasets
Numerous studies have investigated the development of datasets or simulation platforms to create interactive RL environments. RecoGym (Rohde et al., 2018) focuses on simulation bandit environments under e-commerce recommendation setting, while Virtual-Taobao (Shi et al., 2019) provides a user simulator trained on historical behavior data. RecSim (Ie et al., 2019) is a simulation platform supporting sequential user interactions and easy configuration of environments. Additionally, SOFA (Huang et al., 2020) is the first simulator that accounts for interaction biases for optimization and evaluation. In the latest work, RL4RS (Wang et al., 2023) provides a validated simulation environment, advanced evaluation methods, and a real dataset. KuaiSim (Zhao et al., 2023) offers an environment that provides user feedback with multi-behavior, supporting three levels of recommendation tasks.
However, most of these studies merely provide an interactive environment, lacking integrated experimental procedures and other crucial components in RL-based RSs, such as diverse policies and state trackers. RL4RS is the most similar work to our library, with the following distinctions: 1) They focus on the aspects of datasets and evaluation, whereas our goal is to offer an easy-to-use framework for developing new models and facilitating experiments conduction. 2) They support only specific slate recommendation scenarios, whereas our library caters to broader and more commonly used scenarios.
@@ 2.2.2. Frameworks and Libraries
It is worth noting that these libraries primarily serve bandit-based RSs and do not incorporate RL policies. Moreover, some well-established libraries designed for classic RL scenarios(e.g. rllib (Liang et al., 2018), tianshou (Weng et al., 2022)) cannot be directly applied to recommendations due to the lack of recommendation environment construction, state modeling, and so on.
@@ 4.1. Library Overview
The overall architecture of EasyRL4Rec is illustrated in Figure 1. The library is structured around four core modules: Environment, Policy, StateTracker, and Collector, each addressing distinct stages of the reinforcement learning interaction process. These modules build upon the foundations provided by the Gymnasium and Tianshou libraries. The Buffer serves as a fundamental data structure for organizing raw data trajectories, while the Trainer and Evaluator act as executors, managing the entire process. In the following, we will delve into the designs of these core modules. The details of the training and evaluation pipeline can be found in Section 5.
@@ 4.4. StateTracker
Figure 2. Visualzation of data/trajectories stored in Buffer. To support simultaneous interactions in multiple environments, Buffer comprises interaction data from $n$ environments, with different trajectories in each environment represented by distinct colors and the presence of the start symbol.
@@ 5.2. Evaluation
Furthermore, EasyRL4Rec provides abundant evaluation metrics that are commonly used in the field of Reinforcement Learning and Recommender Systems, which are summarised in Table 3. To evaluate the long-term effects of RL policies, we introduce Cumulative Reward ($\text{R}_{\text{cumu}}$) $\sum_{t}r_{t}$ to measure cumulative gain and for one interaction trajectory, which is usually applied in the RL scenario. In addition, Interaction Length ($Length$) and Average Reward ($\text{R}_{\text{avg}}$ measure the length of the interaction trajectory and the single-round reward, respectively. We also provide commonly used metrics in traditional RSs for evaluating user models or other recommendation models, such as Normalized Discounted Cumulative Gain (NDCG (Järvelin and Kekäläinen, 2002)), HitRate, etc.
@@ 7. Experiments
Table 4. Performance comparison between model-free RL methods trained with a user model on three datasets. Policies using continuous action are indicated by the symbol (C).
@@ 7.1.3. Experimental Details
[…]
@@ 7.3. Performance of Batch RL
Table 5. Performance comparison between batch RL methods trained from offline logs.
@@ 7.4. Coverage, Diversity & Novelty
Table 6. Performance of Coverage, Diversity & Novelty on KuaiRec Dataset.
@@ 7.6. Impact of StateTrackers
Table 7. Performance comparison between StateTrackers on the KuaiRec dataset.
@@ 7.7. Impact of Construction Methods
Table 8. Performance comparison between different construction methods on the KuaiRec dataset.
@@ 8. Conclusion & Future Work
In this work, we introduce EasyRL4Rec, an easy-to-use code library specifically crafted for RL-based RSs, simplifying the development and evaluation of RL models. EasyRL4Rec constructs lightweight RL environments based on five public datasets. It offers a unified training and evaluation pipeline, evaluating models from the perspective of long-term benefits. Moreover, our framework facilitates customizable state modeling and action representation, addressing challenges in applying RL to recommender systems. Through comprehensive insightful experiments, we compare the performance of existing models and uncover insights like the Preference Overestimation issue.

### v2

# EasyRL4Rec: An Easy-to-use Library for Reinforcement Learning Based Recommender Systems

@@ Abstract.
Reinforcement Learning (RL)-Based Recommender Systems (RSs) have gained rising attention for their potential to enhance long-term user engagement. However, research in this field faces challenges, including the lack of user-friendly frameworks, inconsistent evaluation metrics, and difficulties in reproducing existing studies. To tackle these issues, we introduce EasyRL4Rec, an easy-to-use code library designed specifically for RL-based RSs. This library provides lightweight and diverse RL environments based on five public datasets and includes core modules with rich options, simplifying model development. It provides unified evaluation standards focusing on long-term outcomes and offers tailored designs for state modeling and action representation for recommendation scenarios. Furthermore, we share our findings from insightful experiments with current methods. EasyRL4Rec seeks to facilitate the model development and experimental process in the domain of RL-based RSs. The library is available for public use [^1] https://github.com/chongminggao/EasyRL4Rec.
Recommender Systems; Reinforcement Learning; Code Library
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657868isbn: 979-8-4007-0431-4/24/07ccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems (RSs) are increasingly becoming integral to various domains, such as e-commerce, social media, and online streaming services. Traditional RSs usually rely on supervised learning methods like collaborative filtering (Schafer et al., 2007) to learn and predict users’ interests. However, these methods often fail to capture the long-term effects, leading to potential issues like Feedback Loops (Mansoury et al., 2020), Filter Bubbles (Flaxman et al., 2016), and other undesirable biases (Chen et al., 2023, 2021).
Reinforcement Learning (RL)-Based Recommender Systems have gained rising attention due to their ability to optimize long-term user engagement. In this setting, the recommendation process is perceived as a multi-step decision-making problem. An agent (recommender) interacts with the environment (users) and receives feedback on actions (items). The main goal of RL-based RSs is to learn an optimal recommendation policy that maximizes cumulative rewards (users’ long-term engagement) through trial and error. Currently, RL-based RSs have been extensively applied in scenarios such as short video recommendations (Cai et al., 2023b; Liu et al., 2023; Cai et al., 2023a; Xue et al., 2023), e-commerce recommendations (Rohde et al., 2018; Zhao et al., 2018a), and beyond.
@@
With increasing attention to RL-based RSs, research in this field is confronted with the following three challenges.
  • Absence of an easy-to-use framework. Currently, there is a lack of easy-to-use frameworks for academic research. Research for RL-based RSs requires interactive environment building. However, existing resources neither use mass datasets nor simulated data generated by simulators, making them inconvenient for usage. In addition, it is difficult to directly apply RL libraries (e.g. rllib (Liang et al., 2018)) to recommendation scenarios due to the absence of suitable environments and state modeling.
To tackle the above issues, we implement a comprehensive code library for RL-based RSs, named EasyRL4Rec. It provides an easy-to-use framework with core modules with rich choices and a unified training and evaluating process, aiming to simplify the model development and experimental process in the domain of RL-based RS. The library is composed of four core modules: Environment, Policy, StateTracker, and Collector, which cater to different stages of the RL interaction process. Environment, built from lightweight static datasets, provides feedback on upcoming actions. The Policy module selects the optimal action based on the current state, which is encoded by StateTracker. Collector bridges the interactions between Environment and Policy. To address the difficulty of obtaining user states from environments, we implement multiple StateTrackers for state modeling, encompassing today’s popular methods in sequential modeling (Hidasi et al., 2015; Tang and Wang, 2018; Kang and McAuley, 2018; Yuan et al., 2019). Since items in recommendation systems are discrete, EasyRL4Rec includes a mechanism to convert continuous actions to discrete items, allowing for continuous action-based policies.
Moreover, EasyRL4Rec provides a unified training and evaluation procedure by Trainer and Evaluator executors. Evaluator supports three modes, allowing for the removal of recommended items and a quit mechanism. The Trainer module offers two training paradigms: learning directly from offline logs or training with a pre-trained user model. With this unified framework, we conduct comprehensive experiments on classic RL models and several recent work (Xin et al., 2020; Gao et al., 2023) and present insightful results. The main contributions of this work can be summarized as follows:
  • Easy-to-use Framework. EasyRL4Rec provides an easy-to-use framework for RL-based RSs. We construct lightweight RL environments based on five public datasets encompassing diverse domains, which are easy to follow for researchers. Moreover, the design of core modules with rich options reduces the complexity of developing a new model.
@@
  • Tailored Designs for Recommendation Scenarios. In response to challenges when applying RL algorithms in practical recommender systems, we have developed customizable modules for state modeling and action representation, with a conversion mechanism to support continuous action-based policies.
  • Insightful Experiments for RL-based RSs. With EasyRL4Rec, we conduct comprehensive experiments to compare the performance of classic RL models and some recent work. We present insightful experimental results from various perspectives.
@@ 2.1. Development of RL-Based RSs
Reinforcement Learning (RL), a branch of machine learning, focuses on agents learning the ability of decision-making through environmental feedback [^2] RL in this work specifically refers to deep reinforcement learning.. Recently, RL-Based Recommender Systems (RSs) have gained considerable attention due to their ability to model recommendation as a multi-step decision-making process and enhance the long-term benefits.
Numerous studies have investigated the application of RL in recommender systems.  Shani et al. (2005) first formulated the recommendation process as a Markov Decision Process (MDP) and utilized model-based RL methods. Zhao et al. (2018b) adapt a DQN (Mnih et al., 2013) architecture to incorporate positive and negative feedback from users.  Zheng et al. (2018) applies the Dueling DQN algorithm to news recommendation. Chen et al. (2019, 2022) extend REINFORCE (Williams, 1992) and off-policy actor-critic algorithm to recommendation.  Xin et al. (2020) proposed to utilize self-supervision signals to empower RL-based RSs. Unlike most methods based on discrete actions,  Cai et al. (2023a); Liu et al. (2023); Xue et al. (2023) investigate the applications of continuous action-based RL and improvements of long-term user engagement in short-video scenarios. In addition,  Ren et al. (2023) focus on effective state representations, while recent work (Gao et al., 2022c, 2023) addresses specific issues when applying RL in RSs.
@@ 2.2. Resources for RL-Based RSs
Existing resources for RL-based RSs can be categorized into two groups: Simulators & Datasets, and Frameworks & Libraries. In Table  1, we summarize the characteristics of existing resources and our EasyRL4Rec regarding modules and experimental procedures.
@@ 2.2.1. Simulators and Datasets
Numerous studies have focused on the development of datasets or simulation platforms to create interactive RL environments. RecoGym (Rohde et al., 2018) focuses on simulation bandit environments under e-commerce recommendation setting, while Virtual-Taobao (Shi et al., 2019) provides a user simulator trained on historical behavior data. RecSim (Ie et al., 2019) is a simulation platform supporting sequential user interactions and easy configuration of environments. Additionally, SOFA (Huang et al., 2020) is the first simulator that accounts for interaction biases for optimization and evaluation. In the latest work, RL4RS (Wang et al., 2023) provides a validated simulation environment, advanced evaluation methods, and a real dataset. KuaiSim (Zhao et al., 2023) offers an environment with multi-behavior feedback, supporting three levels of recommendation tasks.
However, most of these studies merely provide an interactive environment, lacking integrated experimental procedures and other crucial components in RL-based RSs, such as diverse policies and state trackers. RL4RS is the most similar work to our library, with the following distinctions: 1) They focus on datasets and evaluation, whereas our goal is to offer an easy-to-use framework for developing new models and facilitating experiments. 2) They support only specific slate recommendation scenarios, whereas our library caters to broader and more commonly used scenarios.
@@ 2.2.2. Frameworks and Libraries
These libraries primarily serve bandit-based RSs and do not incorporate RL policies. Moreover, some well-established libraries designed for classic RL scenarios(e.g. rllib (Liang et al., 2018), tianshou (Weng et al., 2022)) cannot be directly applied to recommendations due to the lack of recommendation environment construction and state modeling.
@@ 4.1. Library Overview
The overall architecture of EasyRL4Rec is illustrated in Figure 1. The library is composed of four core modules: Environment, Policy, StateTracker, and Collector, each addressing distinct stages of the reinforcement learning interaction process. These modules build upon the foundations provided by the Gymnasium and Tianshou (Weng et al., 2022) libraries. The Buffer serves as a fundamental data structure for organizing raw data trajectories, while the Trainer and Evaluator act as executors, managing the entire process. In the following, we will delve into the designs of these core modules. The details of the training and evaluation pipeline can be found in Section 5.
@@ 4.4. StateTracker
Figure 2. Visualization of data/trajectories stored in Buffer. To support simultaneous interactions in multiple environments, Buffer comprises interaction data from $n$ environments, with different trajectories in each environment represented by distinct colors and the presence of the start symbol.
@@ 5.2. Evaluation
Furthermore, EasyRL4Rec provides abundant evaluation metrics that are commonly used in the field of Reinforcement Learning and Recommender Systems, which are summarised in Table 3. To evaluate the long-term effects of RL policies, we introduce Cumulative Reward ($\text{R}_{\text{cumu}}$) $\sum_{t}r_{t}$ to measure the cumulative gain of one interaction trajectory, which is usually applied in the RL scenario. In addition, Interaction Length ($Length$) and Average Reward ($\text{R}_{\text{avg}}$ measure the length of the interaction trajectory and the single-round reward, respectively. We also provide commonly used metrics in traditional RSs for evaluating user models or other recommendation models, such as Normalized Discounted Cumulative Gain (NDCG (Järvelin and Kekäläinen, 2002)), HitRate, etc.
@@ 7. Experiments
Table 4. Performance comparison between model-free RL methods trained with a user model on three datasets. Policies using continuous action are indicated by the symbol (C). The meaning of underlining and bold should be included in caption. The best results are highlighted in bold, and the second-best results are underlined.
@@ 7.1.3. Experimental Details
In the training stage, all policies are trained with 100 epochs, with the default learning rate 1e-3. We apply the first training paradigm, i.e. learning from offline logs, to train batch RL algorithms. When training model-free RL algorithms, we pre-train a user model (DeepFM (Guo et al., 2017)) to guide the online learning of RL policies. For the evaluation process, after each epoch, the policy undergoes evaluation using 100 episodes (i.e., interaction trajectories), and the maximum recommended sequence length is limited to 30. We set the random seed to 2023 for consistency and reproducibility. The code for replicating all experimental results is available in our library.
@@ 7.3. Performance of Batch RL
Table 5. Performance comparison between batch RL methods trained from offline logs. The best results are highlighted in bold, and the second-best results are underlined.
@@ 7.4. Coverage, Diversity & Novelty
Table 6. Performance of Coverage, Diversity & Novelty on the KuaiRec Dataset. The best results are highlighted in bold, and the second-best results are underlined.
@@ 7.6. Impact of StateTrackers
Table 7. Performance comparison between StateTrackers on KuaiRec. The best results are highlighted in bold, and the second-best results are underlined.
@@ 7.7. Impact of Construction Methods
Table 8. Performance comparison between different construction methods on KuaiRec. The best results are highlighted in bold, and the second-best results are underlined.
@@ 8. Conclusion & Future Work
In this work, we introduce EasyRL4Rec, an easy-to-use code library specifically crafted for RL-based RSs, simplifying the development and evaluation of RL models. EasyRL4Rec constructs lightweight RL environments based on five public datasets. It offers a unified training and evaluation pipeline, evaluating models from the perspective of long-term benefits. Moreover, EasyRL4Rec facilitates customizable state modeling and action representation, addressing challenges in applying RL to recommender systems. Through comprehensive experiments, we compare the performance of existing models and share our findings from various perspectives.